Circular No. 5906

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) J. V. Scotti, Lunar and Planetary Laboratory, reports his measure-

ments of ten of the nuclei of this comet from images obtained (under difficult conditions in the dawn) by T. Gehrels and himself with the 0.9-m Spacewatch telescope at Kitt Peak. The measurements are of the nine nuclei discussed on IAUC 5892-5893 and what is possibly nucleus 20 of Jewitt et al. (1993, Bull. Am. Astron. Soc. 25, 1042). The positions below refer to the brightest nucleus (nucleus 7). Positions for the other nuclei are given on MPEC 1993-X04.

1993 UT  α_{2000}  δ_{2000}
Dec. 9.53912  14^{h}06^{m}16.^{s}83  −13^{o}40′53′′.5
9.54373  14 06 17.09  −13 40 54.2
14.54727  14 09 54.91  −13 59 34.0

MPEC 1993-X04 also gives improved orbital elements, by the undersigned, for the nine nuclei previously discussed, those for nucleus 7 being:

Epoch = 1994 May 8.0 TT

T = 1994 Mar. 29.9945 TT  ω = 354.^{o}9892
e = 0.207491  Ω = 220.9258 }2000.0
q = 5.379569 AU  i = 5.8254

a = 6.788020 AU n^{o} = 0.0557300 P = 17.685 years It now appears that the last encounter with Jupiter was to 0.0006 AU

on 1992 July 7.8 UT and that the adjustments to the 1992 prediscovery ephemeris on IAUC 5801 should be: Feb. 8, Δα = −1^{m}.24, Δδ = −12′.2; Mar. 9, −0^{m}.93, −9′.3; Apr. 8, −0^{m}.64, −6′.1. The IAUC 5893 ephemeris should be adjusted as follows to yield the position of nucleus 17: 1993 Dec. 9, Δα = +0^{m}.12, Δδ = +0′.2; 1994 Jan. 8, +0^{m}.15, +0′.4; Feb. 7, +0^{m}.19, +0′.5; Mar. 9, +0^{m}.23, +0′.6; Apr. 8, +0^{m}.28, +0′.9; May 8, +0^{m}.32, +1′.2; June 7, +0^{m}.37, +2′.2; July 7, +0^{m}.51, +3′.3. For nucleus 1 the adjustments are: 1993 Dec. 9, Δα = +0^{m}.02, Δδ = −0′.4; 1994 Jan. 8, +0^{m}.03, −0′.5; Feb. 7, +0^{m}.04, −0′.6; Mar. 9, +0^{m}.05, −0′.8; Apr. 8, +0^{m}.04, −0′.8; May 8, +0^{m}.03, −0′.9; June 7, 0^{m}.00, −0′.8; July 7, −0^{m}.04, −0′.9. The approximate fractional distances along the nuclear train and Jupiter collision dates in July 1994 are: nucleus 17, 0.00, 17.6; 15, 0.15, 18.3; 14, 0.28, 18.8; 12, 0.40, 19.4; 11, 0.51, 19.9; 7, 0.69, 20.8; 6, 0.79, 21.3; 5, 0.86, 21.6; 1, 1.00, 22.3.

1993 December 14 Brian G. Marsden